John has a jar in his office that contains 66 coins. Some are dimes, and the rest are quarters. If the total value of the coins is $10.35, how many of each denomination did John have?
I got D= 41 and Q= 25
Could someone please work the problem out so I can check my answer.
Not sure if I did every step correctly.

Answer by ikleyn(53938)

.

                        When you have an answer to a problem,  in which you are not confident,  you can  CHECK  it.

                        Moreover,  you  MUST  check it  (!)

                        Below I will show you how to do it.


    (1)  Add 41 and 25:  41 + 25 == 66, OBVIOUSLY.    ! Correct !


    (2)  Calculate the total: it is  41*10 + 25*25 cents = 410 + 625 cents = 10.35 cents = $10.35.     ! Correct !


So,  your numbers satisfy the condition.  It means that the problem is solved correctly and your answer is correct.

My congrats  (!)

From my post you learned something that you did not know before.

Make your check  ALWAYS  and for  EVERY  SOLVED  problem - it is a  NECESSARY  part to  EACH  PROBLEM  solution.

---------------

By the way, there is a bunch of lessons in this site on coin word problems. See the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Non-typical coin problems
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.